# Bridging the Modality Gap: Enhancing Channel Prediction with Semantically Aligned LLMs and Knowledge Distillation

### v1

@@ VII Conclusion
In this paper, we proposed a LLM channel prediction method based on semantic prompting and cross-modal alignment, named CSI-ALM. CSI-ALM first obtains the feature representation of CSI through a channel embedding module, and then combines pre-trained word embeddings with CSI features through a cross-modal attention mechanism to obtain the aligned CSI features. We also introduced a semantic prompting method that retrieves CSI features and pre-trained linguistic knowledge based on maximum similarity, further enhancing the utilization of the LLM’s pre-trained knowledge. Extensive experimental evaluations demonstrated that CSI-ALM outperformed existing deep learning-based methods, achieving a significant improvement of over 1 dB in prediction accuracy in TDD communication scenarios. Furthermore, our experiments showed that the proposed model lightweighting approach enabled CSI-ALM-Light, with only 0.34M parameters, to achieve performance comparable to CSI-ALM under limited training data conditions—where all models were trained using only 10% of the original training dataset—and significantly outperformed conventional deep learning methods, which is crucial for real-world deployment with limited training data or when applied to new environments.

### v2

@@ VII Conclusion
In this paper, we proposed a LLM channel prediction method based on semantic prompting and cross-modal alignment, named CSI-ALM. CSI-ALM first obtains the feature representation of CSI through a channel embedding module, and then combines pre-trained word embeddings with CSI features through a cross-modal attention mechanism to obtain the aligned CSI features. We also introduced a semantic prompting method that retrieves CSI features and pre-trained linguistic knowledge based on maximum similarity, further enhancing the utilization of the LLM’s pre-trained knowledge. Extensive experimental results demonstrate that CSI-ALM consistently outperforms state-of-the-art deep learning methods across various communication scenarios, achieving substantial performance gains. Furthermore, our experiments showed that the proposed model lightweighting approach enabled CSI-ALM-Light, with only 0.34M parameters, to achieve performance comparable to CSI-ALM under limited training data conditions—where all models were trained using only 10% of the original training dataset—and significantly outperformed conventional deep learning methods, which is crucial for real-world deployment with limited training data or when applied to new environments.